Please answer the following questions. Choose only one correct answer. 1. Which one is correct statement? (1 point) A. Beta of a security is difference of R to market portfolio R B. Beta of a security is change of R to market portfolio R C. Beta of a security is sensitivity of R to market portfolio R D. Beta of a security is ratio of R to market portfolio R 2. Fluctuations of a stock’s return that are due to market-wide news are common risk, also called? (1 point) A. systematic, undiversifiable, or market risk B. diversifiable risk C. general risk D. independent risk 3. Suppose Kazakhtelecom stock is equally likely to have a 45% return or a -25% return. What are its expected return and volatility? (1 point) A. E [R]= 5%, SD (R)=35% B. E [R]= 10%, SD (R)=35% C. E [R]= 8%, SD (R)=35% D. E [R]= 10%, SD (R)=30% 4. We define the net present value (NPV) of a project or investment as…? (1 point) A. NPV = PV(Benefits) - PV(Costs) B. NPV=P(Profit) - Difference C. NPV=PV(Benefits) + PV(Costs) D. PV(Benefits) /PV(Costs) 5. Suppose you invest $1000 in an account paying 10% interest per year. How much will you have in the account in 7 years? (2 points) A. $1948.72 B. $1756.43 C. $1458.72 D. $1148.72 6. Suppose you invest $10,000 in A stock, and $30,000 in B stock. You expect a return of 10% for A and 16% for B. What is your portfolio’s expected return? (2 points) A. 10% B. 7% C. 14,5% D. 13% 7. How do we call risks that share no correlation? (1point) A. Standard risk B. Common risk C. General risk D. Independent risk 8. Suppose Air Astana’s stock has a beta of 0.50, whereas SCAT’s beta is 1.25. If the risk-free rate is 4%, and the expected return of the market portfolio is 10%, what is the expected return of an equally weighted portfolio of Air Astana and SCAT stocks, according to the CAPM? (2 points) A. 9,25% B. 9% C. 10% D. 7,25% 9. Suppose you estimate that HALYK bank’s stock has a beta of 1.45. A similar process for Forte bank with beta of 0.79. If the risk-free interest rate is 3% and you estimate the market’s expected return to be 8%, calculate the equity cost of capital for HALYK and Forte banks (2 points) A. 10,25%, 6,95% B. 10%, 6% C. 8%, 6.95% D. 5%, 7% 10. You have just graduated with an MBA and have always been a coffee aficionado. During your two years you spent many nights in Peet’s Coffee and Tea (PEET), and are wondering whether you might replicate their success in locations where they have not yet expanded. To develop your financial plan, estimate the cost of capital of this opportunity assuming a risk-free rate of 3% and a market risk premium of 5%. Estimated beta is 0.83. (2 points) A. 10% B. 7.15% C. 9.3% D. 5%
Answer #1

Answer 1. Option C. Beta of a security is sensitivity of R to market portfolio R

Answer 2. Option A. systematic, undiversifiable, or market risk

Answer 3. Expected return = 0.5*45+(0.5*-25) = 10%  

Standard Deviation = [(0.45-0.1)^2*0.5+(-0.25-0.1)^2*0.5]^0.5 = 35%

Option B. E [R]= 10%, SD (R)=35%  

Answer 4. Option A. NPV = PV(Benefits) - PV(Costs).

Answer 5. Amount = 1000*(1+10%)^7 = 1948.72 Option A. $1948.72

Answer 6. return = (10%*10000+16%*30000) / 40000 = 14.5% Option C. 14,5%  

Answer 7.  Option D. Independent risk

Answer 8. Return = 0.5*(4%+0.5*(10%-4%)+ 0.5*(4%+1.25*(10%-4%))=9.25% Option  A. 9,25%

Answer 9. Cost of equity HALYK bank’s stock=3%+1.45(8%-3%)= .10.25%

Cost of equity Forte bank =3%+0.79(8%-3%) =6.95%

Option A. 10,25%, 6,95%

Answer 10. the cost of capital =3%+0.83*5%= 7.15% Option B. 7.15%